Consider the linear equation
Y ′(x) = λY (x) + (1 − λ) cos(x) − (1 + λ) sin(x), quadY (0) = 1

. The true solution is Y (x) = sin(x) + cos(x). Solve this problem using Euler’s method with several values of λ and h, for 0 ≤ x ≤ 10. Comment on the results.

(a) λ = −1; h = 0.5, 0.25, 0.125 (b) λ = 1; h = 0.5,0.25,0.125

(c) λ = −5; h = 0.5, 0.25, 0.125, 0.0625 (d) λ=5;h=0.0625
Answer #1

(a) del, -1h=0.5 y(a) = = -1 YOU +(-1) COS(W) + -(1+61)) Biu(3) > Ya)=-461) +2 Costa) Yo1 Y=1, to=o. foc.Y) = - YOU) + 2 CUa) % Euler Method clear all clc; format short for k=1:3 lambda=-1 h=1/24k f=@(X,Y)lambda*Y+(1-lambda)*cos(x)- (1+lambda)*sin(x_Y_TrueSol = 0 1.0000 1.0000 0.5000 1.5000 1.3570 1.0000 1.6276 1.3818 1.5000 1.3541 1.0682 2.0000 0.7478 0.4932 2.5000 -0.01.5000 1.2011 1.0682 1.7500 0.9362 0.8057 2.0000 0.6130 0.4932 2.2500 0.2517 0.1499 2.5000 -0.1253 -0.2027 2.7500 -0.4946 -0.b) % Euler Method clear all clc; format short for k=1:3 lambda=1 h=1/24k f=@(X,Y)lambda*Y+(1-lambda)*cos(x)- (1+lambda)*sin(x0.5000 1.5000 1.3570 1.0000 1.7706 1.3818 1.5000 1.8144 1.0682 2.0000 1.7241 0.4932 2.5000 1.6768 -0.2027 3.0000 1.9168 -0.842.0000 1.2464 0.4932 2.2500 1.1033 0.1499 2.5000 0.9901 -0.2027 2.7500 0.9384 -0.5426 3.0000 0.9822 -0.8489 3.2500 1.1572 -1.h = 0.1250 X_Y_TrueSol = 0 1.0000 1.0000 0.1250 1.1250 1.1169 0.2500 1.2345 1.2163 0.3750 1.3269 1.2968 0.5000 1.4012 1.35704.0000 1.2562 -1.4104 4.1250 1.6024 -1.3866 4.2500 2.0108 -1.3411 4.3750 2.4859 -1.2746 4.5000 3.0325 -1.1883 4.6250 3.6560 -8.5000 182.6340 0.1965 8.6250 205.2636 0.0203 8.7500 230.7422 -0.1561 8.8750 259.4288 -0.3301 9.0000 291.7268 -0.4990 9.1250

Here I have done for 1st two cases.

Thank you so much sir